OpenStudy (korosh23):

Pre-cal 12 Question! Polynomial equation. Determine an equation of a polynomial of degree 4 that has -1/2 as a root of multiplicity 3, and 2x^2-x-1as a factor.

OpenStudy (korosh23):

p(x)=a(x+1/2)^3 (2x^2-x-1) p(x)= a (x+1/2)^3 (x-1)(2x+1) How is it possible as a degree of 4? This would be degree of 5.

zepdrix (zepdrix):

Try to examine the (2x+1) factor. What root does that lead to? :)

OpenStudy (korosh23):

x= -1/2

zepdrix (zepdrix):

Good good good. Which means you only need `two more` to get multiplicity 3, ya?

OpenStudy (korosh23):

why multiplicity of 3?

zepdrix (zepdrix):

Because the question is insisting that it have exactly multiplicity 3 at that root, yes? :o

OpenStudy (korosh23):

right :) sorry

OpenStudy (korosh23):

so (x+1/2)^3

zepdrix (zepdrix):

You had originally thought of something like this:\[\large\rm p(x)= a\left(x+\frac{1}{2}\right)^3 (x-1)\left(2x+1\right)\]Let's try something interesting. Let's factor a 2 out of each term in the last bracket.\[\large\rm p(x)= 2a\left(x+\frac{1}{2}\right)^3 (x-1)\left(x+\frac{1}{2}\right)\] You are correct that this is a quintic equation, which is too much! So it turns out we have too many (x+1/2) powers, yes? Because the second degree polynomial that they gave us contained `one of those roots`.

zepdrix (zepdrix):

Ignore the 2 I pulled out, just absorb it into the `a` out front.\[\large\rm p(x)= a\left(x+\frac{1}{2}\right)^{\color{red}{3}} (x-1)\left(x+\frac{1}{2}\right)\]Do you see how you can fix your polynomial?

OpenStudy (korosh23):

I believe it is p(x)=a(x+1/2)^2 (x−1) (x+1/2)

zepdrix (zepdrix):

Ah that looks better! \c:/

OpenStudy (korosh23):

but how to take out that extra degree? @zepdrix

OpenStudy (lochana):

yes. try x = -1/2 for 2x^2 - x -1. you will get zero again. that means (2x -1) also a factor of 2x^2 - x -1

OpenStudy (korosh23):

@lochana yes you are right

zepdrix (zepdrix):

I think you're misunderstanding the question :) It's not saying `root -1/2 multiplicity 3` AND SEPARATELY `2x^2-x-1` No no no ^ it's not saying that. There is overlap in this information. It's saying `root -1/2 multiplicity 3` some of which might come from this `2x^2-x-1`

OpenStudy (korosh23):

oh I see what you mean. I have to adjust the degree of the factors to achieve the polynomial degree of 4. I get it know. :) Thank you!

zepdrix (zepdrix):

ya you didn't have to do anything fancy to fix it, just use your eraser ^^ hehe

OpenStudy (korosh23):

lol